OpenStudy (anonymous):

using L’Hôpital’s Rule evaluate the limit 1)lim 1nx^2/x^2-1 x→1 2)lim e^x/2/x x→ ∞

OpenStudy (anonymous):

@BenjaminEE

OpenStudy (anonymous):

In problem one. What is the first part? 1n x^2? Is that n * x^2?

OpenStudy (anonymous):

And the second one: is it: \[\frac{ e^{x/2} }{ x }\]

OpenStudy (anonymous):

and the other one is #36

OpenStudy (anonymous):

Aha sorry, one sec :)

OpenStudy (anonymous):

lol its okay (:

OpenStudy (anonymous):

Ok the first one: (ln x^2) / (x^2 -1 ) Derive both sides: (1/x^2 * 2x) / (2x) = 1/x^2, input 1 into x and you'll get: 1/1^2 = 1

OpenStudy (anonymous):

And the second one: e^(1/2 x) / x Derive both sides: 1/2 * e^(1/2 x) / 1 = e^(1/2 x) / 2 Now as x grows towards infinity so will e^(1/2 x) / 2 do Which means that: lim (x -> infinity) e^(1/2 x) / x = + infinity

OpenStudy (anonymous):

yay !i actually got the second part right

OpenStudy (anonymous):

thank you ! that helped alot :)